Suppose that a category of world class runners are known to run a marathon (26 miles) in an average of 146 minutes with a standard deviation of 15 minutes. Consider 49 of the races. Let X = the average of the 49 races.

Find the probability that the average of the sample will be between 143 and 147 minutes in these 49 marathons. (Round your answer to four decimal places.)

Find the 60th percentile for the average of these 49 marathons. (Round your answer to two decimal places.)

______ min

Find the median of the average running times.

_____min
Answer #1

Solution :

Given that,

mean = \mu = 146

standard deviation = \sigma = 15

n = 49

\mu\bar x = \mu = 146

\sigma\bar x = \sigma / \sqrt n = 15 / \sqrt 49 = 2.14

a) P(143 < \bar x < 147)  

= P[(143 - 146) /2.14 < (\bar x - \mu \bar x) / \sigma \bar x < (147 - 146) / 2.14)]

= P(-1.40 < Z < 0.47)

= P(Z < 0.47) - P(Z < -1.40)

Using z table,  

= 0.6808 - 0.0808

= 0.6000

b) Using standard normal table,

P(Z < z) = 60%

= P(Z < z) = 0.60

= P(Z < 0.25) = 0.60

z = 0.25

Using z-score formula  

\bar x = z * \sigma \bar x+ \mu \bar x

\bar x = 0.25 * 2.14 + 146

\bar x = 146.54 min.

c) Using standard normal table,

P(Z < z) = 50%

= P(Z < z) = 0.50

= P(Z < 0) = 0.50

z = 0

Using z-score formula  

\bar x = z * \sigma \bar x+ \mu \bar x

\bar x = 0 * 2.14 + 146

\bar x = 146

median = 146 min.